\begin{thebibliography}{81}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Balunovic et~al.(2022)Balunovic, Dimitrov, Jovanovi{\'c}, and Vechev}]{balunovic2022lamp}
Mislav Balunovic, Dimitar Dimitrov, Nikola Jovanovi{\'c}, and Martin Vechev. 2022.
\newblock Lamp: Extracting text from gradients with language model priors.
\newblock \emph{Advances in Neural Information Processing Systems}, 35:7641--7654.

\bibitem[{Bao et~al.(2023)Bao, Zhang, Zhang, Wang, Feng, and He}]{bao2023tallrec}
Keqin Bao, Jizhi Zhang, Yang Zhang, Wenjie Wang, Fuli Feng, and Xiangnan He. 2023.
\newblock Tallrec: An effective and efficient tuning framework to align large language model with recommendation.
\newblock In \emph{Proceedings of the 17th ACM Conference on Recommender Systems}, pages 1007--1014.

\bibitem[{Berlioz et~al.(2015)Berlioz, Friedman, Kaafar, Boreli, and Berkovsky}]{berlioz2015applying}
Arnaud Berlioz, Arik Friedman, Mohamed~Ali Kaafar, Roksana Boreli, and Shlomo Berkovsky. 2015.
\newblock Applying differential privacy to matrix factorization.
\newblock In \emph{Proceedings of the 9th ACM Conference on Recommender Systems}, pages 107--114.

\bibitem[{Bourtoule et~al.(2021)Bourtoule, Chandrasekaran, Choquette-Choo, Jia, Travers, Zhang, Lie, and Papernot}]{bourtoule2021machine}
Lucas Bourtoule, Varun Chandrasekaran, Christopher~A Choquette-Choo, Hengrui Jia, Adelin Travers, Baiwu Zhang, David Lie, and Nicolas Papernot. 2021.
\newblock Machine unlearning.
\newblock In \emph{2021 IEEE Symposium on Security and Privacy (SP)}, pages 141--159. IEEE.

\bibitem[{Carranza et~al.(2023)Carranza, Farahani, Ponomareva, Kurakin, Jagielski, and Nasr}]{carranza2023privacy}
Aldo~Gael Carranza, Rezsa Farahani, Natalia Ponomareva, Alex Kurakin, Matthew Jagielski, and Milad Nasr. 2023.
\newblock Privacy-preserving recommender systems with synthetic query generation using differentially private large language models.
\newblock \emph{arXiv preprint arXiv:2305.05973}.

\bibitem[{Chang et~al.(2010)Chang, Thompson, Wang, and Yao}]{chang2010towards}
Chih-Cheng Chang, Brian Thompson, Hui Wang, and Danfeng Yao. 2010.
\newblock Towards publishing recommendation data with predictive anonymization.
\newblock In \emph{Proceedings of the 5th ACM Symposium on Information, Computer and Communications Security}, pages 24--35.

\bibitem[{Chen et~al.(2020)Chen, Li, Wu, Hong, Wang, and Zhou}]{chen2020secure}
Chaochao Chen, Liang Li, Bingzhe Wu, Cheng Hong, Li~Wang, and Jun Zhou. 2020.
\newblock Secure social recommendation based on secret sharing.
\newblock \emph{arXiv preprint arXiv:2002.02088}.

\bibitem[{Chen et~al.(2022)Chen, Sun, Zhang, and Ding}]{chen2022recommendation}
Chong Chen, Fei Sun, Min Zhang, and Bolin Ding. 2022.
\newblock Recommendation unlearning.
\newblock In \emph{Proceedings of the ACM Web Conference 2022}, pages 2768--2777.

\bibitem[{Chen et~al.(2021)Chen, Liu, Chen, Peng, and Huang}]{chen2021secrec}
Jinrong Chen, Lin Liu, Rongmao Chen, Wei Peng, and Xinyi Huang. 2021.
\newblock Secrec: A privacy-preserving method for the context-aware recommendation system.
\newblock \emph{IEEE Transactions on Dependable and Secure Computing}, 19(5):3168--3182.

\bibitem[{Deng et~al.(2021)Deng, Wang, Li, Shang, Liu, Rajasekaran, and Ding}]{deng2021tag}
Jieren Deng, Yijue Wang, Ji~Li, Chao Shang, Hang Liu, Sanguthevar Rajasekaran, and Caiwen Ding. 2021.
\newblock Tag: Gradient attack on transformer-based language models.
\newblock \emph{arXiv preprint arXiv:2103.06819}.

\bibitem[{Ding et~al.(2023)Ding, Qin, Yang, Wei, Yang, Su, Hu, Chen, Chan, Chen et~al.}]{ding2023parameter}
Ning Ding, Yujia Qin, Guang Yang, Fuchao Wei, Zonghan Yang, Yusheng Su, Shengding Hu, Yulin Chen, Chi-Min Chan, Weize Chen, et~al. 2023.
\newblock Parameter-efficient fine-tuning of large-scale pre-trained language models.
\newblock \emph{Nature Machine Intelligence}, 5(3):220--235.

\bibitem[{Erkin et~al.(2012)Erkin, Veugen, Toft, and Lagendijk}]{erkin2012generating}
Zekeriya Erkin, Thijs Veugen, Tomas Toft, and Reginald~L Lagendijk. 2012.
\newblock Generating private recommendations efficiently using homomorphic encryption and data packing.
\newblock \emph{IEEE transactions on information forensics and security}, 7(3):1053--1066.

\bibitem[{Fan et~al.(2022)Fan, Zhao, Chen, Su, Gao, Wang, Liu, Wang, Xu, Chen et~al.}]{fan2022comprehensive}
Wenqi Fan, Xiangyu Zhao, Xiao Chen, Jingran Su, Jingtong Gao, Lin Wang, Qidong Liu, Yiqi Wang, Han Xu, Lei Chen, et~al. 2022.
\newblock A comprehensive survey on trustworthy recommender systems.
\newblock \emph{arXiv preprint arXiv:2209.10117}.

\bibitem[{Ganju et~al.(2018)Ganju, Wang, Yang, Gunter, and Borisov}]{ganju2018property}
Karan Ganju, Qi~Wang, Wei Yang, Carl~A Gunter, and Nikita Borisov. 2018.
\newblock Property inference attacks on fully connected neural networks using permutation invariant representations.
\newblock In \emph{Proceedings of the 2018 ACM SIGSAC conference on computer and communications security}, pages 619--633.

\bibitem[{Ge et~al.(2022)Ge, Liu, Fu, Tan, Li, Xu, Li, Xian, and Zhang}]{ge2022survey}
Yingqiang Ge, Shuchang Liu, Zuohui Fu, Juntao Tan, Zelong Li, Shuyuan Xu, Yunqi Li, Yikun Xian, and Yongfeng Zhang. 2022.
\newblock A survey on trustworthy recommender systems.
\newblock \emph{ACM Transactions on Recommender Systems}.

\bibitem[{Han et~al.(2024)Han, Gao, Liu, Zhang et~al.}]{han2024parameter}
Zeyu Han, Chao Gao, Jinyang Liu, Sai~Qian Zhang, et~al. 2024.
\newblock Parameter-efficient fine-tuning for large models: A comprehensive survey.
\newblock \emph{arXiv preprint arXiv:2403.14608}.

\bibitem[{Harte et~al.(2023)Harte, Zorgdrager, Louridas, Katsifodimos, Jannach, and Fragkoulis}]{harte2023leveraging}
Jesse Harte, Wouter Zorgdrager, Panos Louridas, Asterios Katsifodimos, Dietmar Jannach, and Marios Fragkoulis. 2023.
\newblock Leveraging large language models for sequential recommendation.
\newblock In \emph{Proceedings of the 17th ACM Conference on Recommender Systems}, pages 1096--1102.

\bibitem[{Hsieh et~al.(2023)Hsieh, Li, YEH, Nakhost, Fujii, Ratner, Krishna, Lee, and Pfister}]{hsieh2023distilling}
Cheng-Yu Hsieh, Chun-Liang Li, CHIH-KUAN YEH, Hootan Nakhost, Yasuhisa Fujii, Alex~Jason Ratner, Ranjay Krishna, Chen-Yu Lee, and Tomas Pfister. 2023.
\newblock Distilling step-by-step! outperforming larger language models with less training data and smaller model sizes.
\newblock In \emph{The 61st Annual Meeting Of The Association For Computational Linguistics}.

\bibitem[{Hu et~al.(2021)Hu, Wallis, Allen-Zhu, Li, Wang, Wang, Chen et~al.}]{hu2021lora}
Edward~J Hu, Phillip Wallis, Zeyuan Allen-Zhu, Yuanzhi Li, Shean Wang, Lu~Wang, Weizhu Chen, et~al. 2021.
\newblock Lora: Low-rank adaptation of large language models.
\newblock In \emph{International Conference on Learning Representations}.

\bibitem[{Hu et~al.(2023)Hu, Wang, Lan, Xu, Lim, Bing, Xu, Poria, and Lee}]{hu2023llm}
Zhiqiang Hu, Lei Wang, Yihuai Lan, Wanyu Xu, Ee-Peng Lim, Lidong Bing, Xing Xu, Soujanya Poria, and Roy Ka-Wei Lee. 2023.
\newblock Llm-adapters: An adapter family for parameter-efficient fine-tuning of large language models.
\newblock In \emph{The 2023 Conference on Empirical Methods in Natural Language Processing}.

\bibitem[{Hu et~al.(2024)Hu, Zhang, Xiao, Wang, Feng, and He}]{hu2024exact}
Zhiyu Hu, Yang Zhang, Minghao Xiao, Wenjie Wang, Fuli Feng, and Xiangnan He. 2024.
\newblock Exact and efficient unlearning for large language model-based recommendation.
\newblock \emph{arXiv preprint arXiv:2404.10327}.

\bibitem[{Huang et~al.(2019)Huang, Liu, and Tang}]{huang2019privacy}
Weiming Huang, Baisong Liu, and Hao Tang. 2019.
\newblock Privacy protection for recommendation system: a survey.
\newblock In \emph{Journal of Physics: Conference Series}, volume 1325, page 012087. IOP Publishing.

\bibitem[{Illman and Temple(2019)}]{illman2019california}
Erin Illman and Paul Temple. 2019.
\newblock California consumer privacy act.
\newblock \emph{The Business Lawyer}, 75(1):1637--1646.

\bibitem[{Jumonji et~al.(2021)Jumonji, Sakai, Sun, and Ku}]{jumonji2021privacy}
Seiya Jumonji, Kazuya Sakai, Min-Te Sun, and Wei-Shinn Ku. 2021.
\newblock Privacy-preserving collaborative filtering using fully homomorphic encryption.
\newblock \emph{IEEE Transactions on Knowledge and Data Engineering}, 35(3):2961--2974.

\bibitem[{Kaur et~al.(2019)Kaur, Kumar, and Obaidat}]{kaur2019multi}
Harmanjeet Kaur, Neeraj Kumar, and Mohammad~S Obaidat. 2019.
\newblock Multi-party secure collaborative filtering for recommendation generation.
\newblock In \emph{2019 IEEE Global Communications Conference (GLOBECOM)}, pages 1--6. IEEE.

\bibitem[{Kim et~al.(2018)Kim, Koo, Kim, Yoon, Shin, and Kim}]{kim2018efficient}
Jinsu Kim, Dongyoung Koo, Yuna Kim, Hyunsoo Yoon, Junbum Shin, and Sungwook Kim. 2018.
\newblock Efficient privacy-preserving matrix factorization for recommendation via fully homomorphic encryption.
\newblock \emph{ACM Transactions on Privacy and Security (TOPS)}, 21(4):1--30.

\bibitem[{Leszczynski et~al.(2023)Leszczynski, Zhang, Ganti, Balog, Radlinski, Pereira, and Chaganty}]{leszczynski2023talk}
Megan Leszczynski, Shu Zhang, Ravi Ganti, Krisztian Balog, Filip Radlinski, Fernando Pereira, and Arun~Tejasvi Chaganty. 2023.
\newblock Talk the walk: Synthetic data generation for conversational music recommendation.
\newblock \emph{arXiv preprint arXiv:2301.11489}.

\bibitem[{Li et~al.(2023{\natexlab{a}})Li, Zhang, Liu, and Chen}]{li2023large}
Lei Li, Yongfeng Zhang, Dugang Liu, and Li~Chen. 2023{\natexlab{a}}.
\newblock Large language models for generative recommendation: A survey and visionary discussions.
\newblock \emph{arXiv preprint arXiv:2309.01157}.

\bibitem[{Li et~al.(2006)Li, Li, and Venkatasubramanian}]{li2006t}
Ninghui Li, Tiancheng Li, and Suresh Venkatasubramanian. 2006.
\newblock t-closeness: Privacy beyond k-anonymity and l-diversity.
\newblock In \emph{2007 IEEE 23rd international conference on data engineering}, pages 106--115. IEEE.

\bibitem[{Li et~al.(2024)Li, Chen, Zheng, Liu, and Wang}]{li2024making}
Yuyuan Li, Chaochao Chen, Xiaolin Zheng, Junlin Liu, and Jun Wang. 2024.
\newblock Making recommender systems forget: Learning and unlearning for erasable recommendation.
\newblock \emph{Knowledge-Based Systems}, 283:111124.

\bibitem[{Li et~al.(2023{\natexlab{b}})Li, Chen, Zheng, Zhang, Gong, Wang, and Chen}]{li2023selective}
Yuyuan Li, Chaochao Chen, Xiaolin Zheng, Yizhao Zhang, Biao Gong, Jun Wang, and Linxun Chen. 2023{\natexlab{b}}.
\newblock Selective and collaborative influence function for efficient recommendation unlearning.
\newblock \emph{Expert Systems with Applications}, 234:121025.

\bibitem[{Lin et~al.(2023{\natexlab{a}})Lin, Tang, Tang, Yang, Dang, and Han}]{lin2023awq}
Ji~Lin, Jiaming Tang, Haotian Tang, Shang Yang, Xingyu Dang, and Song Han. 2023{\natexlab{a}}.
\newblock Awq: Activation-aware weight quantization for llm compression and acceleration.
\newblock \emph{arXiv preprint arXiv:2306.00978}.

\bibitem[{Lin et~al.(2023{\natexlab{b}})Lin, Chen, Wang, Xi, Qu, Dai, Zhang, Tang, Yu, and Zhang}]{lin2023clickprompt}
Jianghao Lin, Bo~Chen, Hangyu Wang, Yunjia Xi, Yanru Qu, Xinyi Dai, Kangning Zhang, Ruiming Tang, Yong Yu, and Weinan Zhang. 2023{\natexlab{b}}.
\newblock Clickprompt: Ctr models are strong prompt generators for adapting language models to ctr prediction.
\newblock \emph{arXiv preprint arXiv:2310.09234}.

\bibitem[{Lin et~al.(2023{\natexlab{c}})Lin, Dai, Xi, Liu, Chen, Li, Zhu, Guo, Yu, Tang et~al.}]{lin2023can}
Jianghao Lin, Xinyi Dai, Yunjia Xi, Weiwen Liu, Bo~Chen, Xiangyang Li, Chenxu Zhu, Huifeng Guo, Yong Yu, Ruiming Tang, et~al. 2023{\natexlab{c}}.
\newblock How can recommender systems benefit from large language models: A survey.
\newblock \emph{arXiv preprint arXiv:2306.05817}.

\bibitem[{Lin et~al.(2022)Lin, Pan, Yang, and Ming}]{lin2022generic}
Zhaohao Lin, Weike Pan, Qiang Yang, and Zhong Ming. 2022.
\newblock A generic federated recommendation framework via fake marks and secret sharing.
\newblock \emph{ACM Transactions on Information Systems}, 41(2):1--37.

\bibitem[{Liu et~al.(2023{\natexlab{a}})Liu, Liu, Zhou, Lv, Zhou, and Zhang}]{liu2023chatgpt}
Junling Liu, Chao Liu, Peilin Zhou, Renjie Lv, Kang Zhou, and Yan Zhang. 2023{\natexlab{a}}.
\newblock Is chatgpt a good recommender? a preliminary study.
\newblock \emph{arXiv preprint arXiv:2304.10149}.

\bibitem[{Liu et~al.(2023{\natexlab{b}})Liu, Deng, Li, Wang, Zhang, Liu, Wang, Zheng, and Liu}]{liu2023prompt}
Yi~Liu, Gelei Deng, Yuekang Li, Kailong Wang, Tianwei Zhang, Yepang Liu, Haoyu Wang, Yan Zheng, and Yang Liu. 2023{\natexlab{b}}.
\newblock Prompt injection attack against llm-integrated applications.
\newblock \emph{arXiv preprint arXiv:2306.05499}.

\bibitem[{Liu et~al.(2023{\natexlab{c}})Liu, Zhang, Chen, Zhang, Bai, Feng, Cui, Li, and Che}]{liu2023conversational}
Yuanxing Liu, Wei-Nan Zhang, Yifan Chen, Yuchi Zhang, Haopeng Bai, Fan Feng, Hengbin Cui, Yongbin Li, and Wanxiang Che. 2023{\natexlab{c}}.
\newblock Conversational recommender system and large language model are made for each other in e-commerce pre-sales dialogue.
\newblock \emph{arXiv preprint arXiv:2310.14626}.

\bibitem[{Luo et~al.(2023{\natexlab{a}})Luo, He, Zhao, Huang, Zhou, Li, Xiao, Zhan, and Song}]{luo2023recranker}
Sichun Luo, Bowei He, Haohan Zhao, Yinya Huang, Aojun Zhou, Zongpeng Li, Yuanzhang Xiao, Mingjie Zhan, and Linqi Song. 2023{\natexlab{a}}.
\newblock Recranker: Instruction tuning large language model as ranker for top-k recommendation.
\newblock \emph{arXiv preprint arXiv:2312.16018}.

\bibitem[{Luo et~al.(2023{\natexlab{b}})Luo, Ma, Xiao, and Song}]{luo2023improving}
Sichun Luo, Chen Ma, Yuanzhang Xiao, and Linqi Song. 2023{\natexlab{b}}.
\newblock Improving long-tail item recommendation with graph augmentation.
\newblock In \emph{Proceedings of the 32nd ACM International Conference on Information and Knowledge Management}, pages 1707--1716.

\bibitem[{Luo et~al.(2024{\natexlab{a}})Luo, Wang, Zhou, Ma, and Song}]{luo2024large}
Sichun Luo, Jiansheng Wang, Aojun Zhou, Li~Ma, and Linqi Song. 2024{\natexlab{a}}.
\newblock Large language models augmented rating prediction in recommender system.
\newblock In \emph{ICASSP 2024-2024 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP)}, pages 7960--7964. IEEE.

\bibitem[{Luo et~al.(2022{\natexlab{a}})Luo, Xiao, and Song}]{luo2022personalized}
Sichun Luo, Yuanzhang Xiao, and Linqi Song. 2022{\natexlab{a}}.
\newblock Personalized federated recommendation via joint representation learning, user clustering, and model adaptation.
\newblock In \emph{Proceedings of the 31st ACM international conference on information \& knowledge management}, pages 4289--4293.

\bibitem[{Luo et~al.(2024{\natexlab{b}})Luo, Xiao, Zhang, Liu, Ding, and Song}]{10.1145/3664927}
Sichun Luo, Yuanzhang Xiao, Xinyi Zhang, Yang Liu, Wenbo Ding, and Linqi Song. 2024{\natexlab{b}}.
\newblock \href {https://doi.org/10.1145/3664927} {Perfedrec++: Enhancing personalized federated recommendation with self-supervised pre-training}.
\newblock \emph{ACM Trans. Intell. Syst. Technol.}
\newblock Just Accepted.

\bibitem[{Luo et~al.(2024{\natexlab{c}})Luo, Yao, He, Huang, Zhou, Zhang, Xiao, Zhan, and Song}]{luo2024integrating}
Sichun Luo, Yuxuan Yao, Bowei He, Yinya Huang, Aojun Zhou, Xinyi Zhang, Yuanzhang Xiao, Mingjie Zhan, and Linqi Song. 2024{\natexlab{c}}.
\newblock Integrating large language models into recommendation via mutual augmentation and adaptive aggregation.
\newblock \emph{arXiv preprint arXiv:2401.13870}.

\bibitem[{Luo et~al.(2024{\natexlab{d}})Luo, Yao, Zhao, and Song}]{luo2024language}
Sichun Luo, Yuxuan Yao, Haohan Zhao, and Linqi Song. 2024{\natexlab{d}}.
\newblock A language model-based fine-grained address resolution framework in uav delivery system.
\newblock \emph{IEEE Journal of Selected Topics in Signal Processing}.

\bibitem[{Luo et~al.(2022{\natexlab{b}})Luo, Zhang, Xiao, and Song}]{luo2022hysage}
Sichun Luo, Xinyi Zhang, Yuanzhang Xiao, and Linqi Song. 2022{\natexlab{b}}.
\newblock Hysage: A hybrid static and adaptive graph embedding network for context-drifting recommendations.
\newblock In \emph{Proceedings of the 31st ACM International Conference on Information \& Knowledge Management}, pages 1389--1398.

\bibitem[{Luo et~al.(2013)Luo, Chen, and Li}]{luo2013distributed}
Zhifeng Luo, Shuhong Chen, and Yutian Li. 2013.
\newblock A distributed anonymization scheme for privacy-preserving recommendation systems.
\newblock In \emph{2013 IEEE 4th International Conference on Software Engineering and Service Science}, pages 491--494. IEEE.

\bibitem[{Ma et~al.(2023)Ma, Fang, and Wang}]{ma2023llm}
Xinyin Ma, Gongfan Fang, and Xinchao Wang. 2023.
\newblock Llm-pruner: On the structural pruning of large language models.
\newblock \emph{Advances in neural information processing systems}, 36:21702--21720.

\bibitem[{Mantelero(2013)}]{mantelero2013eu}
Alessandro Mantelero. 2013.
\newblock The eu proposal for a general data protection regulation and the roots of the ‘right to be forgotten’.
\newblock \emph{Computer Law \& Security Review}, 29(3):229--235.

\bibitem[{Meng et~al.(2019)Meng, Wang, Shu, Li, Chen, Liu, and Zhang}]{meng2019towards}
Xuying Meng, Suhang Wang, Kai Shu, Jundong Li, Bo~Chen, Huan Liu, and Yujun Zhang. 2019.
\newblock Towards privacy preserving social recommendation under personalized privacy settings.
\newblock \emph{World Wide Web}, 22:2853--2881.

\bibitem[{Neel and Chang(2023)}]{neel2023privacy}
Seth Neel and Peter Chang. 2023.
\newblock Privacy issues in large language models: A survey.
\newblock \emph{arXiv preprint arXiv:2312.06717}.

\bibitem[{Nguyen et~al.(2022)Nguyen, Huynh, Nguyen, Liew, Yin, and Nguyen}]{nguyen2022survey}
Thanh~Tam Nguyen, Thanh~Trung Huynh, Phi~Le Nguyen, Alan Wee-Chung Liew, Hongzhi Yin, and Quoc Viet~Hung Nguyen. 2022.
\newblock A survey of machine unlearning.
\newblock \emph{arXiv preprint arXiv:2209.02299}.

\bibitem[{Nguyen et~al.(2024)Nguyen, Nguyen, Le~Nguyen, Pham, Doan, and Wong}]{nguyen2024backdoor}
Thuy~Dung Nguyen, Tuan Nguyen, Phi Le~Nguyen, Hieu~H Pham, Khoa~D Doan, and Kok-Seng Wong. 2024.
\newblock Backdoor attacks and defenses in federated learning: Survey, challenges and future research directions.
\newblock \emph{Engineering Applications of Artificial Intelligence}, 127:107166.

\bibitem[{Ogunseyi et~al.(2023)Ogunseyi, Avoussoukpo, and Jiang}]{ogunseyi2023systematic}
Taiwo~Blessing Ogunseyi, Cossi~Blaise Avoussoukpo, and Yiqiang Jiang. 2023.
\newblock A systematic review of privacy techniques in recommendation systems.
\newblock \emph{International Journal of Information Security}, 22(6):1651--1664.

\bibitem[{Pan et~al.(2020)Pan, Zhang, Ji, and Yang}]{pan2020privacy}
Xudong Pan, Mi~Zhang, Shouling Ji, and Min Yang. 2020.
\newblock Privacy risks of general-purpose language models.
\newblock In \emph{2020 IEEE Symposium on Security and Privacy (SP)}, pages 1314--1331. IEEE.

\bibitem[{Polat and Du(2005)}]{polat2005svd}
Huseyin Polat and Wenliang Du. 2005.
\newblock Svd-based collaborative filtering with privacy.
\newblock In \emph{Proceedings of the 2005 ACM symposium on Applied computing}, pages 791--795.

\bibitem[{Rajput et~al.(2024)Rajput, Mehta, Singh, Hulikal~Keshavan, Vu, Heldt, Hong, Tay, Tran, Samost et~al.}]{rajput2024recommender}
Shashank Rajput, Nikhil Mehta, Anima Singh, Raghunandan Hulikal~Keshavan, Trung Vu, Lukasz Heldt, Lichan Hong, Yi~Tay, Vinh Tran, Jonah Samost, et~al. 2024.
\newblock Recommender systems with generative retrieval.
\newblock \emph{Advances in Neural Information Processing Systems}, 36.

\bibitem[{Sakuma and Osame(2017)}]{sakuma2017recommendation}
Jun Sakuma and Tatsuya Osame. 2017.
\newblock Recommendation with k-anonymized ratings.
\newblock \emph{arXiv preprint arXiv:1707.03334}.

\bibitem[{Saleem et~al.(2021)Saleem, Rehmani, Crespi, and Minerva}]{saleem2021parking}
Yasir Saleem, Mubashir~Husain Rehmani, Noel Crespi, and Roberto Minerva. 2021.
\newblock Parking recommender system privacy preservation through anonymization and differential privacy.
\newblock \emph{Engineering Reports}, 3(2):e12297.

\bibitem[{Shen et~al.(2023)Shen, Chen, Backes, Shen, and Zhang}]{shen2023anything}
Xinyue Shen, Zeyuan Chen, Michael Backes, Yun Shen, and Yang Zhang. 2023.
\newblock " do anything now": Characterizing and evaluating in-the-wild jailbreak prompts on large language models.
\newblock \emph{arXiv preprint arXiv:2308.03825}.

\bibitem[{Shin et~al.(2018)Shin, Kim, Shin, and Xiao}]{shin2018privacy}
Hyejin Shin, Sungwook Kim, Junbum Shin, and Xiaokui Xiao. 2018.
\newblock Privacy enhanced matrix factorization for recommendation with local differential privacy.
\newblock \emph{IEEE Transactions on Knowledge and Data Engineering}, 30(9):1770--1782.

\bibitem[{Shmueli and Tassa(2017)}]{shmueli2017secure}
Erez Shmueli and Tamir Tassa. 2017.
\newblock Secure multi-party protocols for item-based collaborative filtering.
\newblock In \emph{Proceedings of the eleventh ACM conference on recommender systems}, pages 89--97.

\bibitem[{Shmueli and Tassa(2020)}]{shmueli2020mediated}
Erez Shmueli and Tamir Tassa. 2020.
\newblock Mediated secure multi-party protocols for collaborative filtering.
\newblock \emph{ACM Transactions on Intelligent Systems and Technology (TIST)}, 11(2):1--25.

\bibitem[{Shu et~al.(2023)Shu, Gu, Zhang, Zhang, Lu, Li, and Gu}]{shu2023rah}
Yubo Shu, Hansu Gu, Peng Zhang, Haonan Zhang, Tun Lu, Dongsheng Li, and Ning Gu. 2023.
\newblock Rah! recsys-assistant-human: A human-central recommendation framework with large language models.
\newblock \emph{arXiv preprint arXiv:2308.09904}.

\bibitem[{Wang et~al.(2024)Wang, Lin, Chen, Yang, Tang, Zhang, and Yu}]{wang2024towards}
Hangyu Wang, Jianghao Lin, Bo~Chen, Yang Yang, Ruiming Tang, Weinan Zhang, and Yong Yu. 2024.
\newblock Towards efficient and effective unlearning of large language models for recommendation.
\newblock \emph{arXiv preprint arXiv:2403.03536}.

\bibitem[{Wang et~al.(2023)Wang, Ren, Zhou, Lu, Luo, Shi, Zhang, Song, Zhan, and Li}]{wang2023mathcoder}
Ke~Wang, Houxing Ren, Aojun Zhou, Zimu Lu, Sichun Luo, Weikang Shi, Renrui Zhang, Linqi Song, Mingjie Zhan, and Hongsheng Li. 2023.
\newblock Mathcoder: Seamless code integration in llms for enhanced mathematical reasoning.
\newblock In \emph{The Twelfth International Conference on Learning Representations}.

\bibitem[{Wei et~al.(2024)Wei, Ren, Tang, Wang, Su, Cheng, Wang, Yin, and Huang}]{wei2024llmrec}
Wei Wei, Xubin Ren, Jiabin Tang, Qinyong Wang, Lixin Su, Suqi Cheng, Junfeng Wang, Dawei Yin, and Chao Huang. 2024.
\newblock Llmrec: Large language models with graph augmentation for recommendation.
\newblock In \emph{Proceedings of the 17th ACM International Conference on Web Search and Data Mining}, pages 806--815.

\bibitem[{Yan et~al.(2024)Yan, Li, Xu, Dong, Zhang, Ren, and Cheng}]{yan2024protecting}
Biwei Yan, Kun Li, Minghui Xu, Yueyan Dong, Yue Zhang, Zhaochun Ren, and Xiuzheng Cheng. 2024.
\newblock On protecting the data privacy of large language models (llms): A survey.
\newblock \emph{arXiv preprint arXiv:2403.05156}.

\bibitem[{Yao et~al.(2024{\natexlab{a}})Yao, Duan, Xu, Cai, Sun, and Zhang}]{yao2024survey}
Yifan Yao, Jinhao Duan, Kaidi Xu, Yuanfang Cai, Zhibo Sun, and Yue Zhang. 2024{\natexlab{a}}.
\newblock A survey on large language model (llm) security and privacy: The good, the bad, and the ugly.
\newblock \emph{High-Confidence Computing}, page 100211.

\bibitem[{Yao et~al.(2024{\natexlab{b}})Yao, Luo, Zhao, Deng, and Song}]{yao2024can}
Yuxuan Yao, Sichun Luo, Haohan Zhao, Guanzhi Deng, and Linqi Song. 2024{\natexlab{b}}.
\newblock Can llm substitute human labeling? a case study of fine-grained chinese address entity recognition dataset for uav delivery.
\newblock In \emph{Companion Proceedings of the ACM on Web Conference 2024}, pages 1099--1102.

\bibitem[{Yue et~al.(2021)Yue, He, Zeng, and McAuley}]{yue2021black}
Zhenrui Yue, Zhankui He, Huimin Zeng, and Julian McAuley. 2021.
\newblock Black-box attacks on sequential recommenders via data-free model extraction.
\newblock In \emph{Proceedings of the 15th ACM Conference on Recommender Systems}, pages 44--54.

\bibitem[{Zeng et~al.(2024)Zeng, Yue, Jiang, and Wang}]{zeng2024GPTFedRec}
Huimin Zeng, Zhenrui Yue, Qian Jiang, and Dong Wang. 2024.
\newblock \href {http://arxiv.org/abs/2403.04256} {Federated recommendation via hybrid retrieval augmented generation}.

\bibitem[{Zhang et~al.(2024{\natexlab{a}})Zhang, Long, Guo, Fang, Song, Liu, Zhou, Zhang, Liu, and Yang}]{zhang2024FedPA}
Chunxu Zhang, Guodong Long, Hongkuan Guo, Xiao Fang, Yang Song, Zhaojie Liu, Guorui Zhou, Zijian Zhang, Yang Liu, and Bo~Yang. 2024{\natexlab{a}}.
\newblock \href {http://arxiv.org/abs/2405.04840} {Federated adaptation for foundation model-based recommendations}.

\bibitem[{Zhang et~al.(2024{\natexlab{b}})Zhang, Liu, Liu, Wu, Guo, and Wang}]{zhang2024stealthy}
Jinghao Zhang, Yuting Liu, Qiang Liu, Shu Wu, Guibing Guo, and Liang Wang. 2024{\natexlab{b}}.
\newblock Stealthy attack on large language model based recommendation.
\newblock \emph{arXiv preprint arXiv:2402.14836}.

\bibitem[{Zhang et~al.(2023)Zhang, Xie, Hou, Zhao, Lin, and Wen}]{zhang2023recommendation}
Junjie Zhang, Ruobing Xie, Yupeng Hou, Wayne~Xin Zhao, Leyu Lin, and Ji-Rong Wen. 2023.
\newblock Recommendation as instruction following: A large language model empowered recommendation approach.
\newblock \emph{arXiv preprint arXiv:2305.07001}.

\bibitem[{Zhang et~al.(2024{\natexlab{c}})Zhang, Cao, Wu, Sun, Shen, and Cheng}]{zhang2024lorec}
Kaike Zhang, Qi~Cao, Yunfan Wu, Fei Sun, Huawei Shen, and Xueqi Cheng. 2024{\natexlab{c}}.
\newblock Lorec: Large language model for robust sequential recommendation against poisoning attacks.
\newblock \emph{arXiv preprint arXiv:2401.17723}.

\bibitem[{Zhang et~al.(2021{\natexlab{a}})Zhang, Ren, Wang, Ren, Chen, Hu, and Zhang}]{zhang2021membership}
Minxing Zhang, Zhaochun Ren, Zihan Wang, Pengjie Ren, Zhunmin Chen, Pengfei Hu, and Yang Zhang. 2021{\natexlab{a}}.
\newblock Membership inference attacks against recommender systems.
\newblock In \emph{Proceedings of the 2021 ACM SIGSAC Conference on Computer and Communications Security}, pages 864--879.

\bibitem[{Zhang et~al.(2021{\natexlab{b}})Zhang, Zhang, Ji, and Wang}]{zhang2021trojaning}
Xinyang Zhang, Zheng Zhang, Shouling Ji, and Ting Wang. 2021{\natexlab{b}}.
\newblock Trojaning language models for fun and profit.
\newblock In \emph{2021 IEEE European Symposium on Security and Privacy (EuroS\&P)}, pages 179--197. IEEE.

\bibitem[{Zhao et~al.(2024{\natexlab{a}})Zhao, Wang, Xu, Ren, Ng, and Chua}]{zhao2024llm4fedrec}
Jujia Zhao, Wenjie Wang, Chen Xu, Zhaochun Ren, See-Kiong Ng, and Tat-Seng Chua. 2024{\natexlab{a}}.
\newblock \href {http://arxiv.org/abs/2402.09959} {Llm-based federated recommendation}.

\bibitem[{Zhao et~al.(2023)Zhao, Zhou, Li, Tang, Wang, Hou, Min, Zhang, Zhang, Dong et~al.}]{zhao2023survey}
Wayne~Xin Zhao, Kun Zhou, Junyi Li, Tianyi Tang, Xiaolei Wang, Yupeng Hou, Yingqian Min, Beichen Zhang, Junjie Zhang, Zican Dong, et~al. 2023.
\newblock A survey of large language models.
\newblock \emph{arXiv preprint arXiv:2303.18223}.

\bibitem[{Zhao et~al.(2024{\natexlab{b}})Zhao, Fan, Li, Liu, Mei, Wang, Wen, Wang, Zhao, Tang et~al.}]{zhao2024recommender}
Zihuai Zhao, Wenqi Fan, Jiatong Li, Yunqing Liu, Xiaowei Mei, Yiqi Wang, Zhen Wen, Fei Wang, Xiangyu Zhao, Jiliang Tang, et~al. 2024{\natexlab{b}}.
\newblock Recommender systems in the era of large language models (llms).
\newblock \emph{IEEE Transactions on Knowledge and Data Engineering}.

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt]{article}

\usepackage[preprint]{acl}

\usepackage{times}
\usepackage{latexsym}

\usepackage[T1]{fontenc}

\usepackage[utf8]{inputenc}

\usepackage{microtype}

\usepackage{inconsolata}

\usepackage{colortbl}
\definecolor{mygray}{gray}{.9}
\definecolor{mypink}{rgb}{.99,.91,.95}
\definecolor{mycyan}{cmyk}{.3,0,0,0}

\usepackage{tikz}
\usepackage[edges]{forest}

\usepackage{graphicx}

\title{Privacy in LLM-based Recommendation: Recent Advances and Future Directions}

\author{Sichun Luo$^1$, Wei Shao$^1$, Yuxuan Yao$^1$, Jian Xu$^2$, Mingyang Liu$^1$, Qintong Li$^3$, \\\textbf{Bowei He$^1$, Maolin Wang$^1$, Guanzhi Deng$^1$, Hanxu Hou$^4$, Xinyi Zhang$^5$, Linqi Song$^1$} \\
  $^1$CityUHK \quad $^2$Tsinghua University \quad $^3$HKU \quad $^4$Dongguan University of Technology \quad $^5$CUEB\\
  \texttt{sichun.luo@my.cityu.edu.hk, linqi.song@cityu.edu.hk} 
  }

\begin{document}
\maketitle
\begin{abstract}
Nowadays, 
large language models (LLMs) have been integrated with conventional recommendation models to improve recommendation performance.
However, while most of the existing works have focused on improving the model performance, the privacy issue has only received comparatively less attention. In this paper, we review recent advancements in privacy within LLM-based recommendation, categorizing them into privacy attacks and protection mechanisms. Additionally, we highlight several challenges and propose future directions for the community to address these critical problems.
\end{abstract}

\section{Introduction}

Recommender systems have been proposed to address the information overload problem \cite{luo2022personalized,luo2022hysage,luo2023improving,10.1145/3664927}. Recently, large language models (LLMs) have shown great ability in natural language understanding, generation, and reasoning \cite{zhao2023survey,luo2024language,wang2023mathcoder,yao2024can}.
They have also been integrated into recommendation systems, achieving significant progress and reshaping the paradigms of recommendation \cite{bao2023tallrec,lin2023can,zhao2024recommender}.

Despite the extensive research on LLM-based recommendation systems, most studies have centered on model accuracy, with less attention dedicated to privacy issues. However, safeguarding privacy is paramount due to the critical nature of private information. 
Thus privacy issues are not peripheral; they are central to the ethical use of these technologies and warrant thorough investigation. 

Several surveys have explored LLM-based recommendation systems \cite{lin2023can,li2023large,zhao2024recommender}, privacy issues in recommendation systems \cite{huang2019privacy,fan2022comprehensive,ogunseyi2023systematic}, and privacy concerns specific to LLMs \cite{pan2020privacy,neel2023privacy,yao2024survey,yan2024protecting}. Nevertheless, there has been no comprehensive study that specifically focus on privacy in LLM-based recommendation systems, leaving a significant gap in the literature.

In this paper, we aim to bridge this gap. We first provide an overview of LLM-based recommendation systems and examine the privacy issues inherent to LLMs and their applications in recommendations. Thereafter, we discuss various privacy attacks and the corresponding privacy protection mechanisms. 
Subsequently, we point out some challenges and future directions.
Our contributions are threefold:

\begin{itemize}
    \item 
    To the best of our knowledge, we are the first to conduct an extensive review of recent advances in privacy issues in LLM-based recommendation systems. 
    \item 
    We systematically introduce privacy in LLM-based recommendations, including privacy definition, attacks, and protection methods.
    \item 
    We highlight current challenges and promising research directions for developing privacy-preserving LLM-based recommender systems.
\end{itemize}

\section{Preliminary}

\subsection{LLM-based Recommendation}

LLM-based recommendation can be classified into several categories. Some works use LLMs as feature encoders \cite{harte2023leveraging, rajput2024recommender}, feature engineering tools \cite{wei2024llmrec, luo2024integrating}, scoring/ranking functions \cite{luo2023recranker, lin2023clickprompt}, and for modeling user interaction \cite{liu2023conversational, leszczynski2023talk}. 
Noteworthy advancements have been made in utilizing LLMs as recommenders, either through in context learning \cite{luo2024large,liu2023chatgpt} or supervised fine-tuning \cite{bao2023tallrec,zhang2023recommendation}.
However, while a significant emphasis is placed on model accuracy, the aspect of privacy remains relatively under-explored, thereby presenting a substantial research gap.

\subsection{Privacy Issues}

\subsubsection{Privacy Issues in LLMs}

Privacy concerns pervade the life cycle of LLMs, including pre-training, fine-tuning, and inference stages. We briefly discuss them as follows.

\paragraph{Pre-training Stage}
LLMs are trained on massive datasets, often sourced from publicly available internet data, which can contain sensitive information such as addresses, emails, and Social Security numbers. As model parameters increase, the risk of memorizing and inadvertently outputting this private information also rises, leading to potential privacy breaches.

\paragraph{Fine-tuning Stage}
Fine-tuning (FT) adapts pre-trained LLMs to specific tasks using smaller, domain-specific datasets from fields like healthcare, law, and finance. These datasets can include sensitive information, such as patient privacy data or corporate trade secrets. Despite regulations, LLMs may still memorize private content, and attacks like Membership Inference and Model Inversion can exploit this to reveal sensitive information.

\paragraph{Inference Stage}
During inference, LLMs generate content based on user inputs, which may contain sensitive or personally identifiable information, leading to privacy leaks. Even with assurances from LLM service providers, complete protection of user data is challenging. LLMs can infer user attributes and preferences from interaction history, even without explicit private information, posing significant privacy concerns.

\subsubsection{Privacy in Recommendation}
According to previous works \cite{huang2019privacy,ge2022survey}, the definition of privacy shares similar ingredients.
This includes sensitive user attributes such as identity, gender, age, and address, which necessitate stringent access restrictions. The concept of ownership dictates that only authorized entities—whether individual users or the platform itself—are permitted to access and control private information. However, these systems face significant threats from malicious entities, both internal and external, who seek to gain unauthorized access or manipulate this information. These adversaries often exploit auxiliary public information to facilitate their infiltration or attacks. Consequently, the primary goal of privacy protection within recommender systems is to uphold the ownership of private information and develop effective countermeasures against such threats. This ensures that users' sensitive data remains secure and their privacy intact, fostering trust in the technology.

\subsubsection{Privacy Issues in LLM-based Recommendation}

Privacy in LLM-based recommendation is the intersection of privacy in LLM and privacy in recommendation, since it may involve the incorporation of these two kinds of models.
For example,
when utilizing LLMs as recommenders, privacy concerns primarily center around the privacy of the LLM itself. Conversely, when employing LLMs as feature augmentors in conjunction with conventional recommender systems, privacy issues are more likely to arise within the recommendation process.


\section{Privacy Attack in LLM-based Recommendation}



Privacy attacks on recommendation include several categories, such as Membership Inference Attacks \cite{zhang2021membership} (i.e., identify whether the target user is utilized to train the target recommendation model),  Property Inference Attacks \cite{ganju2018property} (i.e., infer global sensitive information in the training set), Reconstruction Attacks \cite{meng2019towards} (i.e., using the user’s publicly available attributes and multiple queries to reconstruct their sensitive information), and Model Extraction Attacks \cite{yue2021black} (i.e., extract useful information from model).

There are several kinds of privacy attacks on LLM, such as 
Prompt Hacking \cite{shen2023anything,liu2023prompt}
(i.e. strategically designing and manipulating input prompts so that they can influence LLMs’s output),
Adversarial Attack \cite{zhang2021trojaning,nguyen2024backdoor}
(i.e. manipulating input data to cause the LLM to produce incorrect or unintended outputs),
Gradient Leakage Attack \cite{deng2021tag,balunovic2022lamp}
(i.e. access or infer the gradients or the gradient information, which may obtain access to the model or even compromise its privacy and safety).

Some pioneer works focus on privacy attacks in LLM-based recommendation \cite{zhang2024stealthy,zhang2024lorec}.
\citet{zhang2024stealthy}
reveals that the LLM-based recommendation introduces security vulnerabilities, allowing attackers to subtly manipulate item recommendations without detection,
and further propose a stealthy attack via prompt injection.
Moreover,
\citet{zhang2024lorec} propose
LoRec, a framework that effectively utilizes LLMs to detect and defend against poisoning attacks in sequential recommender systems, significantly improving their robustness.

\section{Privacy Protection in LLM-based Recommendation}
\label{sec:pri_pro}


\subsection{LLM-based Recommendation Unlearning}
\label{sec:mac_unl}

Machine unlearning \cite{bourtoule2021machine} is a privacy-preserving technique aiming at addressing the \textit{"right to be forgotten"} challenge underscored by various privacy regulations \cite{mantelero2013eu,illman2019california}.  Its primary objective is to facilitate the effective and efficient erasure or exclusion of sensitive or unwanted information from the trained machine learning model while maintaining the model performance \cite{nguyen2022survey}. 

Recommendation unlearning, a subset of machine unlearning, is specifically tailored for recommendation scenarios \cite{chen2022recommendation,li2023selective,li2024making}. 
LLM-based recommendation unlearning focuses on the LLM-based recommendation siutation.
E2URec \cite{wang2024towards} enhances unlearning efficiency by updating select LoRA parameters and employs a teacher-student framework to improve the unlearning process.
Besides,
APA \cite{hu2024exact} utilize unique adapters for segmented training data shards, retraining only those adapters impacted by unusable data, and employing parameter-level adapter aggregation with sample-adaptive attention for each test sample.
These methods enable effective and efficient unlearning while preserving recommendation performance.

\subsection{LLM-based Federated Recommendation}
\label{sec:fl}

Recently, some pioneering studies have already equipped the LLM-based recommendation with the paradigm of federated learning to leverage the foundation models while preserving the users' data privacy \cite{zeng2024GPTFedRec,zhang2024FedPA,zhao2024llm4fedrec}. For example, \citet{zhao2024llm4fedrec} adopts a federated split learning manner to selectively retain certain sensitive LLM layers on the client side while offloading non-sensitive layers to the server to reduce the local training burden. Meanwhile, a dynamic balance strategy is also proposed to ensure relatively balanced performance across clients. \citet{zeng2024GPTFedRec} proposes a novel hybrid Retrieval Augmented Generation (RAG) mechanism for extracting the ID-based user patterns and text-based item features, which are then converted into text prompts and fed into the ChatGPT for re-ranking. In this way, generalized features could be extracted from the sparse and heterogeneous data, and the rich knowledge within the pre-trained LLM could be exploited. \citet{zhang2024FedPA} proposes a novel adapter-based foundation model customization for recommendation in a federated setting, where each client trains its local adapter using the private data and the local adapters are further clustered into multiple groups in the server. Finally, the personalized adapter is obtained by combining both the user-level and user-group-level adapters. Then, the pre-trained foundation models equipped with personalized adapters can provide fine-grained recommendation service efficiently.

\subsection{Other Privacy Preserving Techniques}

There are other privacy-preserving techniques have been employed within the broader domain of recommender systems, including non-cryptography-based techniques (e.g., anonymization \cite{luo2013distributed,chang2010towards,sakuma2017recommendation}, randomization \cite{polat2005svd,li2006t}, and differential privacy \cite{berlioz2015applying,shin2018privacy,saleem2021parking}) 
and a myriad of cryptography-based techniques (e.g., homomorphic encryption \cite{erkin2012generating,kim2018efficient,jumonji2021privacy}, secure multiparty computation \cite{shmueli2017secure,kaur2019multi,shmueli2020mediated}, and secret sharing \cite{chen2020secure,chen2021secrec,lin2022generic}). These methods have proven effective in their respective applications, providing robust privacy safeguards. 
Specifically,
RAH \cite{shu2023rah} consists of LLM-based agents and emphasizes user alignment, to address challenges in recommender systems such as accuracy, biases, privacy, and cold-start problems.
\citet{carranza2023privacy} addresses differential privacy in training deep retrieval systems by using DP language models to generate private synthetic queries, improving retrieval quality while maintaining privacy guarantees.


\section{Challenges and Future Directions}

We detail the existing works for privacy in LLM-based recommendation. However, there are also some other potential directions to be explored. In this section, we will illustrate some promising directions for further research on this topic.

\begin{itemize}
    \item

\noindent  \textit{Challenge 1: Lack of Universally Applicability in Privacy-Preserving LLM-based Recommendations.}
The application of LLM-based recommender systems spans a wide array of scenarios, each with unique requirements and specificities. Current research often targets narrow, specific use cases, limiting the broader applicability and adaptability of these methods. This fragmentation hinders the development of universally and comprehensively applicable solutions that can seamlessly integrate across varying contexts and industries.

\textit{Future Directions:}
To address this challenge, there is a pressing need for a unified, general, and comprehensive framework that can accommodate the diverse requirements of different application scenarios while ensuring privacy preservation. This necessitates the development of multi-task, multi-domain solutions that are robust and versatile enough to handle the complexity and variability inherent in real-world applications.

\item 

\noindent \textit{Challenge 2: Efficiency and Effectiveness in Privacy-Preserving LLM-based Recommendations.}
The substantial size of LLMs poses significant challenges in terms of computational efficiency. The sheer scale of these models can lead to resource-intensive operations, making them impractical for many applications where quick, reliable recommendations are crucial. Furthermore, privacy-preserving techniques, such as encryption, often incur additional computational costs. Besides, the trade-off between privacy and accuracy means that some privacy-preserving methods may compromise the model’s effectiveness. Consequently, preserving user privacy effectively and efficiently remains a challenging task.

\textit{Future Directions:}
Innovative approaches are required to enhance the efficiency and effectiveness of LLM-based recommendation systems. Research should focus on optimizing model architectures, leveraging techniques like model pruning \cite{ma2023llm}, quantization \cite{lin2023awq}, and distillation \cite{hsieh2023distilling} to reduce computational overhead without compromising performance. Additionally, exploring methods to improve inference speed and energy efficiency will be crucial for the practical deployment of LLMs in recommender systems.

\item 

\noindent  \textit{Challenge 3: Privacy-Preserving Cloud-Edge Collaboration for LLM-based Recommendation.} Due to the large size of LLM, applying cloud-edge collaborative recommender systems could be a promising direction to relieve the computation and storage burden for edge devices. However, it may not be practical to fine-tune the cloud-side large models without the disclosure of local sensitive data, thus reducing the recommendation performance. Moreover, the tremendous volume of edge users also makes it infeasible to customize the cloud-side model for each individual user. Besides, latency caused by intermediate information transmission and extensive computation overhead highly impacts the user experience.

\textit{Future Directions:} Model compression techniques for LLMs could be a promising direction for deploying recommendation models in the edge devices while parameter-efficient fine-tuning methods \cite{ding2023parameter,han2024parameter}, such as adapter \cite{hu2023llm} and LoRA \cite{hu2021lora}, could also be applied for local model customization, thus avoiding uploading the local data. Considering heterogeneous device capabilities, novel model architectures with multiple exits could also offer a flexible trade-off between the performance and efficiency during the model inference.

\end{itemize}

\section{Conclusion}

In this paper, we provide a comprehensive overview of recent advancements for privacy in LLM-based recommendation. More specifically, we introduce the basic concepts and taxonomy to facilitate a better understanding of this topic. We also summarize the representative methods addressing privacy concerns in LLM-based recommendation systems, covering both privacy attacks and privacy protection. Finally, we analyze potential challenges and suggest possible future research directions in the realm of privacy for LLM-based recommendations.

\bibliography{custom}

\appendix



\end{document}